[figure: 9.7 cm]

METAMAGNETIC BEHAVIOR IN THE Ln(Mn_{1−x}Co_{x})O_{3} (Ln = Nd, Eu, Y) SYSTEMS WITH PEROVSKITE STRUCTURE

I.O. Troyanchuk^{a}, D.D. Khalyavin^{a}, H. Szymczak^{b}, R. Szymczak^{b}, M. Baran^{b}

a Inst. of Solid State and Semiconductor Phys., National Acad. of Sci. of Belarus,

P.Brovki Str. 17, 220072 Minsk, Belarus

b Institute of Physics, Polish Acad. of Sci., Al. Lotników 32/46, 02-668 Warsaw

We report measurements of magnetization in the Ln(Mn_{1−x}Co_{x})O_{3} Ln = Nd, Eu (0 ≤ x ≤ 1) and Y(Mn_{1−x}Co_{x})O_{3} (0.35 ≤ x ≤ 0.5) perovskites as a function of temperature and magnetic field. It is shown that compositions in the range 0.35 ≤ x ≤ 0.8 (Ln = Eu); 0.4 ≤ x ≤ 0.6 (Ln = Nd) and 0.35 ≤ x ≤ 0.45 (Ln = Y) exhibit a metamagnetic behavior associated with a transformation of a magnetic structure from ferrimagnetic (where magnetic moments of Co^{2+} and Mn^{4+} ions are opposite) to ferromagnetic with a parallel orientation of the Co^{2+} and Mn^{4+} magnetic moments. The magnetization behavior depends strongly on oxygen content. Concentrational and H-T phase diagrams are depicted. Neutron diffraction study indicates an ordering of the Co^{2+} and Mn^{4+} ions in the Y(Mn_{0.5}Co_{0.5})O_{3} sample.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : H. Szymczak

Address for correspondence : Prof. dr hab. H. Szymczak Instytut Fizyki PAN Al. Lotników 32/46 02-668 Warszawa

Email address : szymh@ifpan.edu.pl